# Rolling Diffusion Models

### v1

@@ Appendix A Hyperparameters
Table 4: PDE Experiments
@@
Table 5: BAIR Experiment
@@
Table 6: Kinetics-600 Experiment
Appendix B Simulation Details
In this section we present the hyperparameters to generate the Kolmogorov Flow simulation data, see Table 7. It is important to note that to introduce uncertainty into an otherwise deterministic system, we vary the viscosity and density parameters, which must then be inferred from the data. This would also make a standard solver very difficult to use in such a setting. Additionally, due to the chaotic nature of the system, it is not deterministically predictable up to arbitrary precision.
We use the “simple turbulence forcing” option, which combines a driving force with a damping term such that we simulate turbulence in two dimensions.
@@ Appendix B Simulation Details
Appendix C Additional Results
We show in Table 8 the MSE and FSE errors at various time-steps. Note that the MSE model is always optimal in terms of MSE loss, which is as expected. However, in terms of matching the frequency distribution, as measured by FSD, standard diffusion, and in particular rolling diffusion are optimal.
@@ Appendix C Additional Results
Rolling Diffusion (init noise) (2-4) 29.59 / 39.72 47.44 / 287.8 43.39 / 738.4 61.93 / 769.0 214.32 / 735.7 648.53 / 699.1 1238.44 / 670.0
@@
Appendix D Rolling Diffusion Objective
Standard Diffusion
Further, when $T\rightarrow\infty$, taking care that $s\to t$ we get continuous analog of Equation 30 (Kingma et al., 2021):
$\mathcal{L}_{w}$ $:=\frac{1}{2}\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[w(\lambda)\cdot-\frac{d\lambda_{t}}{dt}\lVert\hat{{\bm{{\epsilon}}}}_{\theta}({\bm{z}}_{t};\lambda_{t})-{\bm{{\epsilon}}}\rVert^{2}\right]\,.$ (31)
with $w(\lambda_{t})=1$, where $\lambda_{t}=\log\operatorname{SNR}(t)$. The weighting function $w(\lambda_{t})=1$ is often changed to improve image quality, for instance by being the inverse $-1/\frac{d\lambda_{t}}{dt}$ so that the objective is simply constant over ${\bm{{\epsilon}}}$-loss.
Rolling Diffusion Objective
In rolling diffusion, the signal-to-noise ratio is kept fixed but one has to account for the local time reparameterization. Recall that $t_{k}:=t_{k}(t)$ denotes the local time reparameterization. Concisely we can say $\operatorname{SNR}_{k}(t):=\operatorname{SNR}(t_{k})$. The rolling continuous time objective is:
$\mathcal{L}_{\infty}$ $=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[\sum_{k=1}^{K}-\frac{d\lambda(t_{k})}{dt_{k}}\frac{t_{k}(t)}{dt}\cdot\lVert{\bm{{\epsilon}}}^{k}-\hat{{\bm{{\epsilon}}}}^{k}_{\theta}({\bm{z}}_{{\bm{{\epsilon}}},t};t)\rVert^{2}\right]\,,$ (32)
where $\lambda(t):=\log\operatorname{SNR}(t)$.
Recall the frame categorization of the main paper, i.e.,
$\text{clean}(s,t)$ :=$:=\{k\mid s_{k}=t_{k}=0\}\,,$ (33)
$\text{noise}(s,t)$ :=$:=\{k\mid s_{k}=t_{k}=1\}\,,$ (34)
$\text{win}(s,t)$ :=$:=\{k\mid s_{k}\in[0,1),t_{k}\in(s_{k},1]\}\,,$ (35)
$\mathcal{L}_{\infty}$ $=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[\sum_{k\in\text{win}(t)}-\frac{d\lambda(t_{k})}{dt_{k}}\frac{t_{k}(t)}{dt}\cdot\lVert{\bm{{\epsilon}}}^{k}-\hat{{\bm{{\epsilon}}}}^{k}_{\theta}({\bm{z}}_{{\bm{{\epsilon}}},t};t)\rVert^{2}\right]\,,$ (36)
This derivation shows a subtle difference between the global perspective (as used above) and the local perspective. If variational bounds are used, they are equivalent and the reparametrization of the time derivative takes into account the edge conditions such as fully noisy frames and fully determined frames. However, constant ${\bm{{\epsilon}}}$ or $v$ losses on the global perspective do differ from the local perspective, as they do not vanish to zero at $t=0$ and $t=1$.
Empirically, ${\bm{{\epsilon}}}$-loss on the window achieved good performance against $v$-loss. Our rolling diffusion objective could be viewed in two ways: As a constant ${\bm{{\epsilon}}}$ defined on the window, or a globally defined constant ${\bm{{\epsilon}}}$-loss that masks out everything but the window, essentially taking into account the derivative of the time reparametrization outside of the window.
Appendix E Algorithms
We present algorithmic outlines for training of rolling diffusion as well as sampling at the boundary. The algorithm for autoregressive rollout was presented in the main text.
Algorithm 2 Rolling Diffusion: Training
  repeat
    Sample ${\bm{x}}$ from $\mathcal{D}_{\mathrm{tr}}$, $t\sim U(0,1)$, $y\sim B(\beta)$
    if $y$ then
       Compute local time $t_{w}^{\text{init}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    else
       Compute local time $t_{w}^{\text{lin}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    end if
    Compute $\alpha_{t_{w}}$ and $\sigma_{t_{w}}$ for all $w=0,\dots,W-1$
    Compute $\hat{{\bm{x}}}\leftarrow f_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}};t)$
    Update $\theta$ using $L_{\text{loc},\theta}({\bm{x}};t,{\bm{{\epsilon}}})$
  until Converged
Algorithm 3 Rolling Diffusion: Sampling at the boundary
  ${\bm{z}}_{1}\leftarrow\operatorname{concat}({\bm{x}},{\bm{z}}_{1})$
  for $t=1,(T-1)/T,\dots,1/T$ do
    Compute local times $t_{w}^{\text{init}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    Sample ${\bm{z}}_{t-1/T}\sim p_{\theta}({\bm{z}}_{t-1/T}|{\bm{z}}_{t})$
  end for
  Return ${\bm{z}}_{0}$, which now has local times $(0/W,1/W,\dots,(W-1)/W)$
Appendix F Hyperparameter Search for $\beta$
Task Training regiment cond-gen FVD
stride=8 Rolling $\beta=0.1$ (5-11) 39.8
Rolling $\beta=0.2$ (5-11) 44.1
Rolling $\beta=0.5$ (5-11) 46.3
Rolling $\beta=0.7$ (5-11) 43.0
Rolling $\beta=0.8$ (5-11) 46.0
Rolling $\beta=0.9$ (5-11) 47.5
stride=1 Rolling $\beta=0.7$ (5-11) 216
Rolling $\beta=0.8$ (5-11) 227
Rolling $\beta=0.9$ (5-11) 211
Table 9: Comparison of oversampling rolling v.s. init noise on Kinetics-600 (stride 8 rollout) with 8192 FVD samples, 100 steps per 11 frames.
@@ Appendix G Rescaled Noise Schedule
For our Kinetics-600 experiments, we used a different noise schedule which can sample from complete noise towards a “rolling state”, i.e., at diffusion times $(\frac{1}{W},\frac{2}{W},\ldots\frac{W}{W})$. From there, we can roll out generation using, e.g., the linear rolling sampling schedule $t_{w}^{\text{lin}}$. The reason is that we hypothesized that the noise schedule $t_{w}^{\text{init}}$ uses a $\operatorname{clip}$ operation, which means that will be sampling in $\text{clean}(s,t)$, which is redundant as outlined in the main paper and Appendix D.
Another schedule that starts from complete noise and ends at the rolling state is the following:
@@
Appendix H Example
Figure 6: Overview of the Rolling Diffusion rollout sampling procedure. Note that the sequences are always partially denoised. Specifically, at the end of the denoising loop they are at “local times” $(0/W,\dots,(W-1)/W)$. After this, the window gets shifted.

### v2

@@ Appendix C Hyperparameters
Table 4: Hyperparameters used for the Kolmogorov flow experiment.
@@
Table 5: Hyperparameters used for the BAIR robot pushing experiment.
@@
Table 6: Hyperparameters used for the Kinetics-600 experiments.
Appendix D Runtime Complexity
Rolling diffusion models do not have an inherent runtime advantage over (batch) autoregressive diffusion models, as both have similar parameter sizes. For fair comparison, one can fix a number of evaluations-per-frame for all models. For example, we allow 32 model evaluations per frame. An autoregressive model would sample a sequence frame by frame, each taking 32 evaluations. A batch-autoregressive model samples, e.g., 8 frames jointly, allowing for 256 model evaluations for this subsequence. In rolling diffusion using a window size of 8, the model gets evaluated 4 times before sliding the window. Upon sampling completion, every frame will be sampled using 32 evaluations.
Given this fixed number of model evaluations per frame, we show in this work that using a rolling sampling strategy can benefit time series generation. However, we note that there are some slight inefficiencies during training. Note that in rolling diffusion, some data is always revealed to the model due to the partial noising schedule. This means that if there is a high overlap between the frames, the model can directly copy the global structure from early frames to the later frames, achieving relatively good reconstruction loss. We believe that because of this, the model gets a less strong signal for learning an “image or video prior”, as it can heavily rely on the conditioning signal. Further experimentation to alleviate this issue is required, perhaps by adjusting the rolling noise schedule or fine-tuning a standard diffusion model using the rolling diffusion loss.
Appendix E Simulation Details
The parameters used to generate the Kolmogorov Flow data are shown in Table 7. It is important to note that to introduce uncertainty into an otherwise deterministic system, we vary the viscosity and density parameters, which must then be inferred from the data. This would also make a standard solver very difficult to use in such a setting. Additionally, due to the chaotic nature of the system, it is not deterministically predictable up to arbitrary precision. We use the ‘simple turbulence forcing’ option, which combines a driving force with a damping term such that we simulate turbulence in two dimensions.
@@ Appendix E Simulation Details
Appendix F Additional Results
F.1 Kolmogorov Flow
We show in Table 8 the MSE and FSE errors at various time-steps. Note that the MSE model is always optimal in terms of MSE loss, which is as expected. However, in terms of matching the frequency distribution, as measured by FSD, standard diffusion, and in particular rolling diffusion are optimal. Averaging ensembles of the rolling diffusion model does not improve the FSD score, but does improve the MSE score.
@@ F.1 Kolmogorov Flow
Rolling Diffusion (init noise) (2-2) 63.21 / 45.62 92.58 / 300.3 144.8 / 748.5 239.1 / 799.2 370.5 / 787.3 529.7 / 767.8 1568.7 / 735.3
Rolling Diffusion (init noise) (2-4) 29.59 / 39.72 47.44 / 287.8 43.39 / 738.4 61.93 / 769.0 214.32 / 735.7 648.53 / 699.1 1238.4 / 670.0
@@
Rolling 10-Ensemble (2-8) 481.6 / 36.4 8590 / 214.7 4 $\cdot 10^{4}$ / 429 5 $\cdot 10^{4}$ / 440 5 $\cdot 10^{4}$ / 429 5 $\cdot 10^{4}$ / 428 5 $\cdot 10^{4}$ / 429
@@
Figure 7: Example rollout of a Kolmogorov flow sample. We depict the vertical velocity field. Note how the MSE model’s intensity decreases as we move further from the conditioning frames.
F.2 Kinetics-600
In Table 2 we compare against TECO (Yan et al., 2023) on the Kinetics-600 dataset. TECO uses 20 conditioning frames and generates 80 new frames at a resolution of 128 by 128 using 256 FVD samples.
Method FVD ($\downarrow$)
TECO 799
Rolling Diffusion (Ours) 685
Table 9: In TECO (Yan et al., 2023), samples are generated conditioning on 20 frames and generating 80 new frames at a resolution of 128 by 128 using 256 FVD samples. Different from other parts of the paper because of the low sample count, FVD is measured by matching the conditioning for the reference samples.
Furthermore, the following plot shows MSE deviations from ground-truth on Kinetics-600 data.
Figure 8: This figure shows MSE as a function of frame distance from the starting point on Kinetics-600 on the 20-12 setting with rollouts until frame 80 on resolution 128 $\times$ 128. As the generated frames are further from the initial 20 conditioning frames, the error between the original and the generated samples increases.
@@ Appendix G Rescaled Noise Schedule
For our Kinetics-600 experiments, we used a different noise schedule which can sample from complete noise towards a “rolling state”, i.e., at diffusion times $(\frac{1}{W},\frac{2}{W},\ldots\frac{W}{W})$. From there, we can roll out generation using, e.g., the linear rolling sampling schedule $t_{w}^{\text{lin}}$. The reason is that we hypothesized that the noise schedule $t_{w}^{\text{init}}$ uses a $\operatorname{clip}$ operation, which means that will be sampling in $\text{clean}(s,t)$, which is redundant as outlined in the main paper and Appendix A.
@@
Appendix H Hyperparameter Search for $\beta$
Task Training regiment Cond-Gen FVD
stride=8 Rolling init rescaled $\beta=0.1$ (5-11) 39.8
steps=24 Rolling init rescaled $\beta=0.2$ (5-11) 44.1
Rolling init rescaled $\beta=0.5$ (5-11) 46.3
Rolling init rescaled $\beta=0.7$ (5-11) 43.0
Rolling init rescaled $\beta=0.8$ (5-11) 46.0
Rolling init rescaled $\beta=0.9$ (5-11) 47.5
Rolling init clip $\beta=0.0$ (5-11) 60.3
Rolling init clip $\beta=0.2$ (5-11) 52.0
Rolling init clip $\beta=0.5$ (5-11) 49.0
Rolling init clip $\beta=0.7$ (5-11) 48.9
stride=1 Rolling init rescaled $\beta=0.7$ (5-11) 216
steps=64 Rolling init rescaled $\beta=0.8$ (5-11) 227
Rolling init rescaled $\beta=0.9$ (5-11) 211
Standard Diffusion (15-1) 1460
Standard Diffusion (5-1) 1369
Standard Diffusion (8-8) 157.1
Standard Diffusion (5-8) 142.4
Table 10: Oversampling $t_{w}^{\text{lin}}$ vs. $t_{w}^{\text{init}}$ noise on Kinetics-600 with 8192 FVD samples. We allow 100 denoising steps per 11 frames.